OpenStudy (anonymous):

Find the 50th term of the sequence 5, –2, –9, –16, ...

hartnn (hartnn):

the first term = a1 =5 the common difference = d = 2--5 = -9-(-2) = ?? then n=50 in an = a+(n-1)d = ?